Write the equation of the line passing through (7,4) and (14,12)
Answer by checkley71(8403)

FIRST WE FIND THE SLOPE
(Y2-Y1)/(X2-X1)=(12-4)/(14-7)=8/7 IS THE SLOPE (m).
NOW WE USE ONE SET OF (X,Y) VALUES AND THE SLOPE (m) 8/7 IN THE LINE EQUATION[Y=mX+b] & SOLVE FOR THE Y INTERCEPT (b).
4=9/8*7+b
4=63/8+b
b=4-63/8
b=(32-68)/8
b=-36/8
b=-4.5
THUS THE LINE BEQUATION IS:
Y=9X/8-4.5